# DeepInsert: Early Layer Bypass for Efficient and Performant Multimodal Understanding

### v2

@@ 7 Limitations & Societal Impact
In terms of societal impact, since our work focuses on (pretrained) MLLMs, we suffer from the standard pitfalls associated with such models, including hallucination, equitable access, lack of accountability, and data attribution. At the same time, our proposed method enhances efficiency, potentially lowering computational costs and energy requirements for multimodal model development while maintaining performance, leading to positive societal, economic, and environmental impacts.
Appendix A Additional Experiments: Interpretability
A.1 Vision-Language Alignment in Pretained Unimodal Models
@@
We begin with some formal exposition to characterize the notion of alignment, borrowed from Huh et al. [36]. A representation is a function $\mathbf{f}:\mathcal{X}\to\mathbb{R}^{n}$ that assigns a feature vector to each input in some data domain $\mathcal{X}$. A kernel, $\mathbf{K}:\mathcal{X}\times\mathcal{X}\to\mathbb{R}$, characterizes how its corresponding representation measures distance/similarity. It is defined as $\mathbf{K}(x_{i},x_{j})=\langle\mathbf{f}(x_{i}),\mathbf{f}(x_{j})\rangle,$ where $\langle\cdot,\cdot\rangle$ denotes the inner product, $x_{i},x_{j}\in\mathcal{X}$, and $\mathbf{K}\in\mathcal{K}$. Finally, a kernel-alignment metric, $\mathbf{m}:\mathcal{K}\times\mathcal{K}\to\mathbb{R}$, measures the similarity between two kernels, that is, how similar is the distance measure induced by one representation to the distance measure induced by another. As in Huh et al. [36], we rely on a mutual nearest-neighbor metric that measures the mean intersection of the $k$-nearest neighbor sets induced by two kernels (normalized by $k$).
@@
In essence, we measure the distance between the kernels induced by the representation spaces of language and vision. First, these representations are obtained by passing paired image-caption data through the respective pretrained models. Then, we rely on a mutual $k$-nearest neighbor alignment metric to estimate the alignment between the corresponding feature spaces. We want to understand how alignment varies as a function of model depth, with the intuition that layers capturing “world semantics" are likely to align better, even across modalities. For the LLaVA v1.5-7b model, its starting components include the vision encoder, clip-vit-large-patch14-336 [67], and the decoder-only LLM vicuna-v1.5-7b model [91]. To compute the alignment metric, we use $k=10$ nearest neighbors over 1024 samples from WIT (Wikipedia-based Image Text; [73]); additional details can be found in Huh et al. [36].
@@
The alignment plot in Fig. 4 is in line with our expectations: the penultimate layer of the vision model aligns more with the latter layers of the LLM. This may point to some shared semantics in the deeper layers of the respective models, and can perhaps be exploited in designing alignment frameworks for VLMs. Note that this alignment is more general. Fig. 5 shows a similar trend in DINO-based encoders [8]. This is remarkable as it points to a more general preexisting alignment, since DINO vision encoders are—unlike CLIP—trained in a self-supervised language-agnostic fashion.
@@
A.2 Where Visual Tokens Matter
@@
Appendix B Layer Selection Heuristic
The key idea is as follows: if there is intrinsic evidence that multimodal tokens are not being used in the early layers, can we skip them even in baseline pretrained models? Our somewhat surprising find is that it is possible to do so, without any finetuning. Specifically, we load the pretrained model weights of the MLLM (LLaVA or LTU) onto the DeepInsert variant, by simply letting the multimodal adapter insert tokens into the latter layer.
Skip Layers (LLaVA) POPE
0 86.0
@@ Appendix B Layer Selection Heuristic
Table 4: Performance across layers, for baseline LLaVA pretrained weights, on POPE benchmark.
Skip Layers (LTU) ESC50 VS
0 81.30 56.39
@@
Table 5: Performance across layers, for baseline LTU pretrained weights, ESC50 and VS audio benchmarks.
From Table 4 and 5, we note that both models retain competency even when layers are skipped. While the performance drops in LLaVA are modest yet noticeable, LTU surprisingly even improves for DI-4, suggesting a strong implicit redundancy, at least for classification benchmarks. We also note that these trends are also indicative of which model/modality are likely to be more susceptible to DeepInsert.
@@
Appendix C Additional Experiments: MLLMs
C.1 BLIP
@@
BLIP (Bootstrapped Language-Image Pretraining) is a pioneering approach for vision-language modeling [44]. It introduces a lightweight Querying Transformer (QFormer) that extracts information from the embeddings of the vision encoder. Vision information is transferred into a set of learnable query tokens (32 embeddings of dimension 768) via the QFormer’s cross-attention layers, which are then fed into the LLM along with the language prompt. Further improvements by: a) scaling the LLM and pretraining dataset [45], and b) multi-task instruction finetuning [15] allows the model to solve a variety of tasks like image captioning, visual question answering, and multimodal reasoning.
@@
In our experiment, we will consider one such model to demonstrate the potential of DeepInsert. Specifically, we consider the VLM’s construction with FlanT5-base [13] (encoder-decoder architecture, 350M parameter model) as our LLM and the Eva-clip-g/14 [75], a CLIP-based vision transformer, for our frozen image encoder. The QFormer follows the default architecture and initialization as in Dai et al. [15]. Since FlanT5-base has an encoder-decoder architecture, insertion bypasses the encoder entirely. Instead of feeding the Qformer outputs (query tokens) as inputs to the encoder, we feed them directly into the decoder by prepending them to the language tokens. The choice of bypassing the encoder is intuitive, since the encoder output space is likely to capture the semantics of the language prompt and therefore align better with the image semantics captured in the QFormer tokens.
@@
Note that we do not intend to follow the training recipe for the InstructBLIP models. Indeed, it is impossible to replicate the model for two reasons: a) lack of information on training recipe, including learning rate schedules and dataset sampling ratios, and b) parts of the LAION dataset, the key component for pretraining, was taken offline so as to comply with U.S. federal law, cf. Thiel [78]. Furthermore, the high-level recipe calls for multi-epoch pretraining on a dataset with roughly 130M image-text pairs, which are beyond the scope of this work. Instead, we define a smaller setup, which still allows us fair testing of DeepInsert but with a reasonable compute budget. Recall that the end-to-end training phase for InstructBLIP involves two steps: the pretraining step with image-caption pairs, followed by the multi-task instruction finetuning step. Consequently, we structure the experiment in a similar vein, by first pretraining efficiency, followed by a comparison post multi-task instruction finetuning.
C.1.1 Pretraining Efficiency
@@
Reflecting the InstructBLIP setup, we consider image-caption pretraining with the COCO captions dataset [11] as the first stage. The goal of this experiment is to compare the level of alignment between the QFormer and the LLM, as training proceeds. Specifically, we measure the best possible captioning score achieved within a fixed number of epochs (1, 10 and 20) for each framework. This is done in accordance with different hyperparameter and learning rate configurations available in the BLIP repository and for each variant, we report the highest performance across configurations. As shown in Fig. 9, we find that our framework not only achieves better peak performance, but also does it in much fewer epochs. We limit our report to 20 epochs as both models achieve over 95% of their respective peak scores within that frame (the peak is achieved at around 50 epochs).
@@
Figure 9: Pretraining comparison for default (QFormer) vs. grounded (Ours) pipelines. Higher score indicates that DeepInsert achieves better alignment, and in much fewer epochs.
C.1.2 Multi-task Training
@@
To emulate InstructBLIP’s training stages, we first pretrain with image-caption pairs for 20 epochs, and then perform multi-task instruction finetuning for captioning and VQA for 15 combined epochs (uniformly sampling from the datasets without replacement). For captioning, we consider the COCO captions dataset as before and for visual question answering, we consider VQAv2 [29] for training. Instruction finetuned VLMs can demontrate remarkable zero-shot abilities on unseen datasets, and so we also consider OKVQA (Outside Knowledge VQA) [58] for zero-shot evaluation. Coco Captions and VQAv2 share the same train-val-test image split that aims to prevent train-test information leakage for a fair evaluation.
Results are given in Table 6. We find that for the designated tasks, our method achieves significant improvements over the standard QFormer. We also see a remarkable gap in zero-shot OKVQA performance, with the grounded model even outperforming the reported performance of some of the largest BLIP-2 models [45]. (InstructBLIP includes OKVQA during training, so no direct comparison is possible.)
Model COCO Captions (BLEU-4) VQAv2 (%) OKVQA (0-shot) (%)
Default (QFormer) 20.9 55.4 28.8
Grounded (Ours) $\bm{36.2}$ $\bm{66.8}$ $\bm{39.0}$
@@
C.2 Training Loss Comparison on LLaVA
@@
Appendix D Efficiency Analysis
@@
$\text{Attention mechanism FLOPs}=4L^{2}d_{\text{model}}+2Ld_{\text{model}}^{2}.$ (2)
@@
$\text{FLOPs per layer for $L$ tokens}=8Ld_{\text{model}}^{2}+4L^{2}d_{\text{model}}+4Ld_{\text{model}}d_{\text{ff}}.$ (4)
@@
$\text{FLOPs}(N_{\text{DI}})=N$ $\Big(8(L_{\text{text}}+L_{\text{mm}})d_{\text{model}}^{2}+4(L_{\text{text}}+L_{\text{mm}})^{2}d_{\text{model}}+4(L_{\text{text}}+L_{\text{mm}})d_{\text{model}}d_{\text{ff}}\Big)$
$-N_{\text{DI}}\Big(8L_{\text{mm}}d_{\text{model}}^{2}+4(2L_{\text{text}}+L_{\text{mm}})L_{\text{mm}}d_{\text{model}}+4L_{\text{mm}}d_{\text{model}}d_{\text{ff}}\Big).$ (5)

### v3

@@ 6 Conclusion & Future Work
Further, since our work focuses on (pretrained) MLLMs, we suffer from the standard pitfalls associated with such models, including hallucination, equitable access, lack of accountability, and data attribution. At the same time, our proposed method enhances efficiency, potentially lowering computational costs and energy requirements for multimodal model development while maintaining performance, leading to positive societal, economic, and environmental impacts.
Appendix A Discussion on Potential Mechanisms Behind Redundancy
Herein, we speculate on possible reasons for early multimodal redundancy. The first hypothesis comes from works on mechanistic interpretability, which suggests that the forward pass of an LLM can be broadly delineated into phases: a) the function determination phase that processes the prompt to determine the task that needs to be executed (for instance, get_capital()) and b) the execution of the function for the given input query (for instance, get_capital(France) = Paris) (Hendel et al., 2023; Lv et al., 2024; Merullo et al., 2024). Based on this, we can question whether a similar delineation holds for LLM-based MLLMs: specifically, if we interpret the multimodal tokens as the input query, then the analogous query execution should take place in the latter layers, which is consistent with our observation that the tokens are used only in the middle layers. Note that query execution here would take an inherently different form, since multimodal tokens are generally either prepended, or inserted in the middle of language prompt.
Another, more theoretically inclined intuition, comes from recent work in associative memories. Associative memory models for understanding retrieval in transformers have gained traction in recent years, in part due to the strong similarity of the respective inner-product attention mechanisms Ramsauer et al. (2021); Smart et al. (2025). Specifically, Betteti et al. (2025) propose a theoretical model for input-driven dynamics for robust retrieval, wherein the retrieval matrix evolves as a function of an additional input, besides the query. Thus, we can extending our previous analogy, by visualizing the instruction part of the prompt as coaxing the model into the space of vision language retrieval. This coaxing takes some initial processing, beyond which the vision query actually comes into effect. Devising a theoretical retrieval model that can model this 2-stage process can help shed light on any potential benefits that arise from these dynamics (along with the observed redundancy).
Appendix B Additional Experiments: Interpretability
B.1 Vision-Language Alignment in Pretained Unimodal Models
@@
We begin with some formal exposition to characterize the notion of alignment, borrowed from Huh et al. (2024). A representation is a function $\mathbf{f}:\mathcal{X}\to\mathbb{R}^{n}$ that assigns a feature vector to each input in some data domain $\mathcal{X}$. A kernel, $\mathbf{K}:\mathcal{X}\times\mathcal{X}\to\mathbb{R}$, characterizes how its corresponding representation measures distance/similarity. It is defined as $\mathbf{K}(x_{i},x_{j})=\langle\mathbf{f}(x_{i}),\mathbf{f}(x_{j})\rangle,$ where $\langle\cdot,\cdot\rangle$ denotes the inner product, $x_{i},x_{j}\in\mathcal{X}$, and $\mathbf{K}\in\mathcal{K}$. Finally, a kernel-alignment metric, $\mathbf{m}:\mathcal{K}\times\mathcal{K}\to\mathbb{R}$, measures the similarity between two kernels, that is, how similar is the distance measure induced by one representation to the distance measure induced by another. As in Huh et al. (2024), we rely on a mutual nearest-neighbor metric that measures the mean intersection of the $k$-nearest neighbor sets induced by two kernels (normalized by $k$).
@@
In essence, we measure the distance between the kernels induced by the representation spaces of language and vision. First, these representations are obtained by passing paired image-caption data through the respective pretrained models. Then, we rely on a mutual $k$-nearest neighbor alignment metric to estimate the alignment between the corresponding feature spaces. We want to understand how alignment varies as a function of model depth, with the intuition that layers capturing “world semantics" are likely to align better, even across modalities. For the LLaVA v1.5-7b model, its starting components include the vision encoder, clip-vit-large-patch14-336 Radford et al. (2021), and the decoder-only LLM vicuna-v1.5-7b model Zheng et al. (2023). To compute the alignment metric, we use $k=10$ nearest neighbors over 1024 samples from WIT (Wikipedia-based Image Text; Srinivasan et al. (2021)); additional details can be found in Huh et al. (2024).
@@
The alignment plot in Fig. 4 is in line with our expectations: the penultimate layer of the vision model aligns more with the latter layers of the LLM. This may point to some shared semantics in the deeper layers of the respective models, and can perhaps be exploited in designing alignment frameworks for VLMs. Note that this alignment is more general. Fig. 5 shows a similar trend in DINO-based encoders Caron et al. (2021). This is remarkable as it points to a more general preexisting alignment, since DINO vision encoders are—unlike CLIP—trained in a self-supervised language-agnostic fashion.
@@
B.2 Where Visual Tokens Matter
@@
Appendix C Layer Selection Heuristic
The key idea is as follows: if there is intrinsic evidence that multimodal tokens are not being used in the early layers, can we skip them even in baseline pretrained models? Our somewhat surprising find is that it is possible to do so, without any finetuning. Specifically, we load the pretrained model weights of the MLLM (LLaVA-7B or LTU-7B) onto the DeepInsert variant, by simply letting the multimodal adapter insert tokens into the latter layer.
Model POPE
LLaVA 86.0
@@ Appendix C Layer Selection Heuristic
(a) LLaVA on POPE
Model ESC50 VS
LTU 81.30 56.39
@@
(b) LTU on ESC50 and VS
Table 5: Performance across layers for baseline pretrained weights on vision (left) and audio (right) benchmarks.
From Table 5 and 5, we note that both models retain competency even when layers are skipped. While the performance drops in LLaVA are modest yet noticeable, LTU surprisingly improves for DI-4, suggesting a strong implicit redundancy, at least for classification benchmarks. We also note that these trends are also indicative of which model/modality are likely to be more susceptible to DeepInsert.
@@
Appendix D Predicting Insert Layer with RL
We consider a pretrained and frozen LLaVA model augmented with a lightweight adapter whose goal is to determine the transformer layer at which multimodal embeddings should be integrated. The adapter takes prompt embeddings as input and outputs a discrete prediction over layers, corresponding to the insertion point for multimodal information.
The adapter is trained using reinforcement learning (RL) with a reward function designed to balance task performance and computational efficiency. Concretely, the reward consists of two components: (i) the negative next-token prediction loss on the given datapoint, which encourages performance retention, and (ii) a redundancy reward that incentivizes skipping later layers. The redundancy reward is normalized to the range $[0,1]$, where 0 indicates no layer skipping and $1$ corresponds to skipping all layers. These two terms are combined via a weighting coefficient $\lambda$, which controls the trade-off between performance and efficiency. By adjusting $\lambda$, the optimization can be biased toward either higher accuracy or reduced computation.
Figure 9: Total reward (red) over optimization steps, decomposed into the performance reward (blue) and the efficiency-related penalty (orange). The dashed horizontal line indicates zero reward.
The adapter is trained on a small subset (10%) of the LLaVA instruction fine-tuning dataset, consisting of samples from COCO, TextVQA, and GQA. The adapter architecture is a simple multilayer perceptron (MLP).
The Figure  9 illustrates the evolution of the reward over time. In the early stages, the total reward exhibits large negative values and high variance, reflecting unstable exploration and frequent violations of efficiency constraints. As training progresses, the policy rapidly improves, with the total reward increasing toward zero and fluctuations becoming less severe. After convergence, the reward remains close to zero for most steps, punctuated by occasional negative spikes. Overall, the trend indicates successful learning of a policy that balances task performance with efficiency considerations, achieving stable behavior while respecting the imposed penalty structure.
The average predicted insertion depth is approximately $0.14$ in normalized layer index, corresponding to layer $4.8$ in a 32-layer transformer (as used in 7B-scale models). This closely aligns with our empirical findings for DI-4.
While further exploration is necessary, these results suggest a principled, data-driven approach for identifying suitable insertion layers for multimodal integration.
Appendix E Integrating with token-reduction techniques
As we mention in Sec. 2, DeepInsert in complementary to token-reduction/token-pruning techniques such as FastV Chen et al. (2024), VTW Lin et al. (2024), and PruMerge Shang et al. (2024), and applying these to DeepInsert models can give further efficiency gains. These techniques drop less important tokens in the attention mechanism, decreasing the computational load of the model. Specifically, we find that DeepInsert models generally respond better to token-pruning methods, compared to the baseline. In line with recommended hyperparameters, for FastV, we use K=3 (layer to start pruning) and R=25% (retention ratio). For DeepInsert, K=3 implies pruning 3 layers after the insertion layer. Similarly, we use K=16 for VTW (implying tokens exit after layer 16). For PruMerge, we use the training free-variant with R=12.5%. Table 6 lists the average performance over AI2D, POPE, and MMMU, when applying the token-reduction techniques, to the baseline LLaVA-7B model as well as the DeepInsert variants, highlighting that the latter is equally or more robust to pruning methods.
Model Baeline Acc. +FastV +VTW +PruMerge
LLaVA 59.4 57.8 (-1.6) 59.9 (+0.5) 59.8(+0.4)
DI-4 59.4 59.2 (-0.2) 59.9 (+0.5) 59.8(+0.4)
DI-8 57.6 58.0 (+0.4) 58.0 (+0.4) 58.1(+0.5)
DI-12 57.3 57.9 (+0.6) 57.9 (+0.6) 57.8(+0.5)
Table 6: Average performance of LLaVA-7B and DeepInsert models, showing that DeepInsert models are equally or more robust to FastV, VTW, and PruMerge.
Appendix F Additional Experiments: MLLMs
F.1 BLIP
@@
BLIP (Bootstrapped Language-Image Pretraining) is a pioneering approach for vision-language modeling Li et al. (2022). It introduces a lightweight Querying Transformer (QFormer) that extracts information from the embeddings of the vision encoder. Vision information is transferred into a set of learnable query tokens (32 embeddings of dimension 768) via the QFormer’s cross-attention layers, which are then fed into the LLM along with the language prompt. Further improvements by: a) scaling the LLM and pretraining dataset Li et al. (2023a), and b) multi-task instruction finetuning Dai et al. (2023) allows the model to solve a variety of tasks like image captioning, visual question answering, and multimodal reasoning.
@@
In our experiment, we will consider one such model to demonstrate the potential of DeepInsert. Specifically, we consider the VLM’s construction with FlanT5-base Chung et al. (2022) (encoder-decoder architecture, 350M parameter model) as our LLM and the Eva-clip-g/14 Sun et al. (2023a), a CLIP-based vision transformer, for our frozen image encoder. The QFormer follows the default architecture and initialization as in Dai et al. (2023). Since FlanT5-base has an encoder-decoder architecture, insertion bypasses the encoder entirely. Instead of feeding the Qformer outputs (query tokens) as inputs to the encoder, we feed them directly into the decoder by prepending them to the language tokens. The choice of bypassing the encoder is intuitive, since the encoder output space is likely to capture the semantics of the language prompt and therefore align better with the image semantics captured in the QFormer tokens.
@@
Note that we do not intend to follow the training recipe for the InstructBLIP models. Indeed, it is impossible to replicate the model for two reasons: a) lack of information on training recipe, including learning rate schedules and dataset sampling ratios, and b) parts of the LAION dataset, the key component for pretraining, was taken offline so as to comply with U.S. federal law, cf. Thiel (2023). Furthermore, the high-level recipe calls for multi-epoch pretraining on a dataset with roughly 130M image-text pairs, which are beyond the scope of this work. Instead, we define a smaller setup, which still allows us fair testing of DeepInsert but with a reasonable compute budget. Recall that the end-to-end training phase for InstructBLIP involves two steps: the pretraining step with image-caption pairs, followed by the multi-task instruction finetuning step. Consequently, we structure the experiment in a similar vein, by first pretraining efficiency, followed by a comparison post multi-task instruction finetuning.
F.1.1 Pretraining Efficiency
@@
Reflecting the InstructBLIP setup, we consider image-caption pretraining with the COCO captions dataset Chen et al. (2015) as the first stage. The goal of this experiment is to compare the level of alignment between the QFormer and the LLM, as training proceeds. Specifically, we measure the best possible captioning score achieved within a fixed number of epochs (1, 10 and 20) for each framework. This is done in accordance with different hyperparameter and learning rate configurations available in the BLIP repository and for each variant, we report the highest performance across configurations. As shown in Fig. 10, we find that our framework not only achieves better peak performance, but also does it in much fewer epochs. We limit our report to 20 epochs as both models achieve over 95% of their respective peak scores within that frame (the peak is achieved at around 50 epochs).
@@
Figure 10: Pretraining comparison for Baseline (QFormer) vs. DeepInsert (Ours) pipelines. Higher score indicates that DeepInsert achieves better alignment, and in much fewer epochs.
F.1.2 Multi-task Training
@@
To emulate InstructBLIP’s training stages, we first pretrain with image-caption pairs for 20 epochs, and then perform multi-task instruction finetuning for captioning and VQA for 15 combined epochs (uniformly sampling from the datasets without replacement). For captioning, we consider the COCO captions dataset as before and for visual question answering, we consider VQAv2 Goyal et al. (2017) for training. Instruction finetuned VLMs can demontrate remarkable zero-shot abilities on unseen datasets, and so we also consider OKVQA (Outside Knowledge VQA) Marino et al. (2019) for zero-shot evaluation. Coco Captions and VQAv2 share the same train-val-test image split that aims to prevent train-test information leakage for a fair evaluation.
Results are given in Table 7. We find that for the designated tasks, our method achieves significant improvements over the standard QFormer. We also see a remarkable gap in zero-shot OKVQA performance, with the grounded model even outperforming the reported performance of some of the largest BLIP-2 models Li et al. (2023a). (InstructBLIP includes OKVQA during training, so no direct comparison is possible.)
Model COCO Cap. (BLEU-4) VQAv2 (%) OKVQA (0-shot) (%)
Baseline 20.9 55.4 28.8
DI (Ours) $\bm{36.2}$ $\bm{66.8}$ $\bm{39.0}$
@@
F.2 Open-LLaVA-NeXT
Herein, we extend DeepInsert to a the newer LLaVA-Next-7B (the open reproduction) Liu et al. (2024b); Chen and Xing (2024), with a much larger vision context yielding performance improvements. Below, we report the average performances in Table 8, indicating the robustness of DeepInsert to exploit early layer redundancy even when number of multimodal tokens are much higher ($\sim$2k). We note here that unlike retraining the baseline models in the main paper, we directly used the checkpoint from Huggingface for the baseline due to limited compute, meaning the DeepInsert models had about 1-2k missing data points, which explains the slight performance drop.
Model AI2D GQA POPE MMMU TextVQA (lite) Avg.
LLaVA-NeXT 64.6 64.2 87.3 39.0 63.4 63.7
DI-4 63.7 63.8 87.4 36.9 64.3 63.2
DI-6 62.7 63.6 87.0 37.1 61.9 62.4
Table 8: Performance of DeepInsert on Open-LLaVA-NeXT-7B.
F.3 Training Loss Comparison on LLaVA
@@
Appendix G Efficiency Analysis
@@
$\text{Attention FLOPs}=4L^{2}d_{\text{model}}+2Ld_{\text{model}}^{2}.$ (2)
@@
FLOPs per layer for $L$ tokens
$=8Ld_{\text{model}}^{2}+4L^{2}d_{\text{model}}+$ $4Ld_{\text{model}}d_{\text{ff}}.$ (4)
@@
$\text{FLOPs}(N_{\text{DI}})$
$=N\Big(8(L_{\text{text}}+L_{\text{mm}})d_{\text{model}}^{2}$
$\qquad\qquad+4(L_{\text{text}}+L_{\text{mm}})^{2}d_{\text{model}}$
$\qquad\qquad\qquad\qquad+4(L_{\text{text}}+L_{\text{mm}})d_{\text{model}}d_{\text{ff}}\Big)$
$-N_{\text{DI}}\Big(8L_{\text{mm}}d_{\text{model}}^{2}$
$\qquad\qquad+4(2L_{\text{text}}+L_{\text{mm}})L_{\text{mm}}d_{\text{model}}$
$\qquad\qquad\qquad\qquad+4L_{\text{mm}}d_{\text{model}}d_{\text{ff}}\Big).$ (5)
@@
Appendix H Potential Risks
Efficiency gains can lower deployment costs, potentially enabling wider, sometimes unguarded use. If models hallucinate or provide ungrounded rationales, this may scale harmful outcomes (e.g., disinformation, low-friction generation of misleading content) when safeguards are absent. As with other MLLMs, the system remains vulnerable to adversarial prompts and distribution shifts. Efficiency itself neither increases nor decreases susceptibility, but wider, cheaper access can increase attack surface (e.g., more attempts, faster iteration by adversaries).
Appendix I AI Usage
AI models (part of Overleaf native tools) were used in improving the writing of the paper.